OpenStudy (anonymous):

Integrate \[ \int 4e^{-x}\cos x \;dx \]

OpenStudy (anonymous):

I'll just keep it simple and do \[ \int e^{-x}\cos x\;dx \]And multiply the \(4\) in later.

OpenStudy (cggurumanjunath):

integrate by parts.

OpenStudy (cggurumanjunath):

use ILATE rule.

OpenStudy (anonymous):

\[ u=\cos x \quad dv=e^{-x}\;dx \]

OpenStudy (cggurumanjunath):